OpenStudy (amigatour):

I'm confused again. I have a given problem: 2K(s)+Cl2(g)--->2KCl(s); find the limiting reactants if K is 1 mole and Cl2 is 1 mole. Write your answer in a chemical formula. My limiting reactant is 0.5mole Cl2. How to write a chemical formula? I did use the coefficient, but it's not working.

OpenStudy (photon336):

You need to multiply each by the molar ratio here's how you can do it What you have 1 mol of Cl2 1 mol of K What you need to ensure that the reaction goes to completion. Remember this isn't a 1:1 molar ratio so one reagent is limiting. 1 mol K x (1Cl2/2K) = 0.5 mol of Cl2 1 mol of Cl2 x (2K/cl2) = 2 mol of K Now what we have 1 mol of each. We need 0.5 mol of Cl2 to get this going and 2 mol of K but we only have 1 mol of K so therefore K is our limiting reagent and we use that to figure out any other calculations ie how much products are produced.

OpenStudy (amigatour):

I see. Thank you. I didn't think the other way. :-/

OpenStudy (photon336):

If you noticed we multiplied by the molar ratio ensuring that what we want is in the numerator and what were multiply the molar ratio by is in the denominator. Ie If we want the number of mol of chlorine we do 1/2 and we want to ensure that potassium is in the denominator. 1.0 mol of K x (Cl2/2K) = 0.5 mol Cl2 Noticed potassium canceled out